# Learning Stackable and Skippable LEGO Bricks for Efficient, Reconfigurable, and Variable-Resolution Diffusion Modeling

### v1

@@ C.2 Ablation Studies: On the effects of LEGO bricks
Figure 8: For a LEGO-Diffusion model consisting of a patch-brick with a patch size of $4\times 4$ and an image-brick with a patch size of $64\times 64$, both with a receptive field size set as $2\times 2$, we visualize how the FID and training time change as the ratio of the depth (number of DiT-based Transformer layers) of the image-brick to that of the patch-brick increases from 0 to 1. Specifically, at the beginning of x-axis, we have a patch-brick with 12 DiT-based Transformer layers and an image-brick with 0 DiT-based Transformer layers, and at the end, we have a patch-brick with 0 DiT-based Transformer layers and an image-brick with 12 DiT-based Transformer layers.
@@
Proportion of layers assigned to the image-brick
We first study the trade-off between the training time and the ratio of the depth (number of DiT block layers) of the image-brick to the overall depth of the LEGO model. For simplicity, we consider a LEGO-Diffusion model consisting of a patch-brick with a patch size of $4\times 4$ and an image-brick with a patch size of $64\times 64$, both with a receptive-field-size set as $2\times 2$. We also let both bricks have the same token embedding dimension, number of attention heads, MLP ratio, etc. Fixing the total depth as 12, we vary the depth assigned to the image-brick from 0 to 12 and show the results in Figure 8. We can observe that when we allocate the majority of layers to the patch-brick, the training is efficient but results in a relatively higher FID. As we assign more layers to the image-brick, the FID improves, but the model requires longer training time. Empirically, we find that when the ratio is greater than 50%, the FID improvement becomes less significant, while efficiency decreases at the same time. This observation motivates us to allocate approximately 50% of the layers to the patch-brick.
@@
To further optimize training efficiency, for the $16\times 16$ patch-brick, we increase its receptive field size from $2\times 2$ to $8\times 8$ and investigate how to assign the six DiT block layers between the $4\times 4$ patch-brick and the $16\times 16$ patch-brick. The results shown on the right panel of Figure 9 reveal that assigning more layers to the $4\times 4$ patch-brick in general results in lower FID but longer training time. Finally, we observe that when the depth of the $16\times 16$ patch-brick is approximately twice that of the $4\times 4$ patch-brick, the model achieves a satisfactory balance between performance and efficiency. These observations have been leveraged in the construction of the LEGO model configurations, which are outlined in Tables 4 and 5. It’s worth noting that additional tuning of configurations, such as employing specific local receptive field sizes, adjusting the embedding dimension of the local receptive field, varying the number of attention heads, and so on for each LEGO brick, holds the potential for further performance improvements. However, we leave these aspects for future exploration and research.
C.3 Leveraging a pretrained diffusion model as a LEGO brick
Table 6: Results of incorporating a pre-trained model as the first brick in LEGO-PR.
Model       VP      VP (frozen) + LEGO-PR   VP (unfrozen) + LEGO-PR
FID 2.5 1.92 1.88
C.4 Additional Generation results
@@
Generating beyond training resolution: Below we first provide additional panorama image generations at resolutions significantly higher than the training images, as shown in Figure 10-12. For all panorama generations (including the top panel in Figure 2), we utilize class-conditional LEGO-PG, trained on ImageNet ($256\times 256$ and $512\times 512$). We first initialize the latent representation of the panorama image with white Gaussian noise at the desired resolution, such as $(1280/8)\times(256/8)\times 4$ for generating $1280\times 256$ RGB images in Figure 10. During the generation process, at each timestep, we employ a sliding stride of 7 and have LEGO predict noise within a window of $(256/8)\times(256/8)$ until the target resolution is covered. The predicted noise at each latent voxel is averaged based on the number of times it is generated, taking into account overlaps. When generating images or patterns with a mixture of class conditions, the idea is to produce diverse outputs that adhere to different class characteristics within a single canvas. This could be achieved by dividing the canvas into different regions, where each region corresponds to a specific class condition.
@@
Figure 11: Paranoma generation of 4K ($3840\times 2160$) resolution with a mixture of class conditions, produced with class-conditional LEGO-PG, trained on ImageNet ($256\times 256$).

### v3

@@ D.2 Ablation Studies: On the effects of LEGO bricks
Proportion of layers assigned to the image-brick
Figure 10: Visualization: how the FID and training time change as the ratio of the depth (number of DiT-based Transformer layers) of the image-brick to that of the patch-brick increases from 0 to 1.
We first study the trade-off between the training time and the ratio of the depth (number of DiT block layers) of the image-brick to the overall depth of the LEGO model. For simplicity, we consider a LEGO-Diffusion model consisting of a patch-brick with a patch size of $4\times 4$ and an image-brick with a patch size of $64\times 64$, both with a receptive-field-size set as $2\times 2$. We also let both bricks have the same token embedding dimension, number of attention heads, MLP ratio, etc. Fixing the total depth as 12, we vary the depth assigned to the image-brick from 0 to 12 and show the results in Figure 10. For a LEGO-Diffusion model consisting of a patch-brick with a patch size of $4\times 4$ and an image-brick with a patch size of $64\times 64$, both with a receptive field size set as $2\times 2$, we visualize how the FID and training time change as the ratio of the depth (number of DiT-based Transformer layers) of the image-brick to that of the patch-brick increases from 0 to 1. Specifically, at the beginning of x-axis, we have a patch-brick with 12 DiT-based Transformer layers and an image-brick with 0 DiT-based Transformer layers, and at the end, we have a patch-brick with 0 DiT-based Transformer layers and an image-brick with 12 DiT-based Transformer layers. We can observe that when we allocate the majority of layers to the patch-brick, the training is efficient but results in a relatively higher FID. As we assign more layers to the image-brick, the FID improves, but the model requires longer training time. Empirically, we find that when the ratio is greater than 50%, the FID improvement becomes less significant, while efficiency decreases at the same time. This observation motivates us to allocate approximately 50% of the layers to the patch-brick.
@@
To further optimize training efficiency, for the $16\times 16$ patch-brick, we increase its receptive field size from $2\times 2$ to $8\times 8$ and investigate how to assign the six DiT block layers between the $4\times 4$ patch-brick and the $16\times 16$ patch-brick. The results shown on the right panel of Figure 11 reveal that assigning more layers to the $4\times 4$ patch-brick in general results in lower FID but longer training time. Finally, we observe that when the depth of the $16\times 16$ patch-brick is approximately twice that of the $4\times 4$ patch-brick, the model achieves a satisfactory balance between performance and efficiency. These observations have been leveraged in the construction of the LEGO model configurations, which are outlined in Tables 5 and 6. It’s worth noting that additional tuning of configurations, such as employing specific local receptive field sizes, adjusting the embedding dimension of the local receptive field, varying the number of attention heads, and so on for each LEGO brick, holds the potential for further performance improvements. However, we leave these aspects for future exploration and research.
D.3 Additional Generation results
@@
Generating beyond training resolution: Below we first provide additional panorama image generations at resolutions significantly higher than the training images, as shown in Figures 12-14. For all panorama generations (including the top panel in Figure 1), we utilize class-conditional LEGO-PG, trained on ImageNet ($256\times 256$ and $512\times 512$). As deployed in Bar-Tal et al. (2023), large content is achievable through a pre-trained text-to-image diffusion model trained on images of lower resolution. In our study, we adopt this methodology to show the efficacy of the LEGO framework in such tasks. Notably, the LEGO framework exhibits enhanced capabilities in managing spatial dependencies, which is a critical aspect in this context. For each paranoma generation, we first initialize the latent representation of the panorama image with white Gaussian noise at the desired resolution, such as $(1280/8)\times(256/8)\times 4$ for generating $1280\times 256$ RGB images in Figure 12. During the generation process, at each timestep, we employ a sliding stride of 7 and have LEGO predict noise within a window of $(256/8)\times(256/8)$ until the target resolution is covered. The predicted noise at each latent voxel is averaged based on the number of times it is generated, taking into account overlaps. When generating images or patterns with a mixture of class conditions, the idea is to produce diverse outputs that adhere to different class characteristics within a single canvas. This could be achieved by dividing the canvas into different regions, where each region corresponds to a specific class condition. In our evaluation, we employed the LPIPS distance metric proposed by Zhang et al. (2018), to assess the similarity between sections of panoramic generations and images from our training set. For each image, we use either a grid-based cropping or a random cropping process. Our analysis involved 100 generated images from various categories. The results are presented in Table 7, including both the average and the standard deviation of the LPIPS distances. Notably, when employing the random cropping method, the quality of panoramas generated using the LEGO model evidently exceeded that of the DiT model. This difference is particularly located in areas comprising outputs from multiple diffusion models, highlighting LEGO’s capacity in managing spatial relationships among image patches.
Table 7: Comparision of large-content generation using LPIPS distance (lower is better).
Crop Grid - 256 Random - 256 Grid - 512 Random - 512
DiT 0.15 $\pm$ 0.07 0.66 $\pm$ 0.13 0.55 $\pm$ 0.07 0.76 $\pm$ 0.06
LEGO-PG 0.14 $\pm$ 0.07 0.21 $\pm$ 0.17 0.36 $\pm$ 0.07 0.37 $\pm$ 0.06
LEGO-PR 0.15 $\pm$ 0.07 0.25 $\pm$ 0.15 0.55 $\pm$ 0.07 0.55 $\pm$ 0.05
@@
Figure 13: Paranoma generation of 4K ($3840\times 2160$) resolution with a mixture of class conditions, produced with class-conditional LEGO-PG, trained on ImageNet ($512\times 512$).